OpenStudy (kaylala):

topic: PROVING IDENTITIES - help me please. i really dont know how to do this (SEE COMMENTS) (trigonometry) 1) [(2+csc x)/(sec x)] - 2 csc x = cot x 2) (1)/(sec x - tan x) = sec x + tan x

OpenStudy (anonymous):

Are you sure about the first equation?

OpenStudy (kaylala):

yep. that's what's given @MIssCiku why? what's wrong with it in your opinion?

OpenStudy (anonymous):

Was that link helpful? and I wrote the first wrongly.. still working it out

OpenStudy (kaylala):

very helpful okay

OpenStudy (loser66):

the second one is easy but the first one is not. There is something wrong with it, please check for the second one , time both numerator and denominator with (sec +tan) , denominator = sec^2 -tan^2 =1 , you have the answer

OpenStudy (anonymous):

Yeah. I thought the first one had a problem. Thanks

OpenStudy (kaylala):

in joining the fractions part how come there're a negative before the whole equation? @MIssCiku

OpenStudy (loser66):

If the last term from the left side is 2 cos , the problem turns true. I mean \[\frac{2+csc x}{secx}-\color{red}{2cosx}=cotx\]

OpenStudy (kaylala):

oh i see @Loser66 if it turns out to be like that ^ how'll it be solved?

OpenStudy (loser66):

\[=\frac{2}{secx}+\frac{cscx}{secx}-2cosx\] \[2cosx +\frac{cosx}{sinx}-2cosx=cot x\]

OpenStudy (kaylala):

@MIssCiku : nevermind my last question "in joining the fractions part how come there're a negative before the whole equation?" i got it already

OpenStudy (kaylala):

@Loser66 how did you get "2cosx+cosxsinx−2cosx=cotx"

OpenStudy (loser66):